OpenStudy (anonymous):

Solve the following system using the substitution method: 3x + y = -3 y = 2x + 7

OpenStudy (astrophysics):

Well it's already set up to use that method for you, so substitute y = 2x+7 into the first equation where y is. 3x(2x+7)=-3 and solve for x, then substitute x back into any of the original equations then solve for y.

OpenStudy (anonymous):

is it inconsistent or dependent

OpenStudy (astrophysics):

I'm not sure, you'll have to do it to figure that out.

OpenStudy (anonymous):

how

OpenStudy (astrophysics):

Well lets solve for x and see

OpenStudy (anonymous):

iss it 1/4

OpenStudy (astrophysics):

\[3x+(2x+7)=-3\]

OpenStudy (astrophysics):

Oh ok, well if it's inconsistent you would get something like 0 = -3, dependent would be -3 = -3

OpenStudy (astrophysics):

This one seems to be dependent.

OpenStudy (anonymous):

ok what about the coordinates

OpenStudy (anonymous):

is there a solution

OpenStudy (astrophysics):

That's what we're trying to solve right now

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

5x + 7=-3?

OpenStudy (astrophysics):

\[3x+2x+7 = -3 \implies 5x=-10 \implies x = -2\] so we have x now, lets solve for y, \[y=2x+7 \implies y=2(-2)+7 \implies y = -4+7 \implies y = 3\] So your coordinates are (-2,3) you can check if I did the algebra right.

OpenStudy (anonymous):

thanks